# arXiv:2602.23381v1[math.GN] 03 Feb 2026

### v1

@@ 2.4 Locally convex spaces
Let $X$ be a locally convex topological vector space and let $\sigma:\mathbb{R}\to\mathbb{R}$ satisfy the univariate universal approximation property of Definition 2.2. Then, for every integer $m\geq 1$ and every compact set $K\subset X$, the TFNN class $\mathcal{N}_{1}^{(m)}(\sigma)$ associated with the feature family $\mathcal{A}(X)=X^{*}$ is dense in $C(K;\mathbb{R}^{m})$ with respect to the uniform norm on $K$. Consequently, for every depth $l\geq 1$, the deep TFNN class $\mathcal{N}_{l}^{(m)}(\sigma)$ is dense in $C(K;\mathbb{R}^{m})$.
@@
Moreover, $\mathcal{E}$ separates points of $K$. Indeed, since $X$ is locally convex, the Hahn–Banach extension theorem holds. It is a consequence of this theorem that for any $x,y\in K$ with $x\neq y$ there exists $f\in X^{*}$ such that $f(x)\neq f(y)$ (see, e.g., [31, Theorem 3.6]). Consequently,
@@
We now apply Theorem 2.1. Let $m\geq 1$, let $K\subset X$ be compact, and let $g\in C(K;\mathbb{R}^{m})$ be arbitrary. Since $\mathcal{A}(X)=X^{*}$ has the $D$-property and $\sigma$ satisfies the univariate universal approximation property, Theorem 2.1 implies that for every $\varepsilon>0$ there exists a TFNN
@@
Finally, since every shallow TFNN can be regarded as a deep TFNN by inserting additional affine and activation layers, we have
$\mathcal{N}_{1}^{(m)}(\sigma)\subset\mathcal{N}_{l}^{(m)}(\sigma)\quad\text{for all }l\geq 1.$
Hence density holds for all depths. ∎
@@
Let $Y$ be a compact metric space and let $V\subset C(Y)$ be compact. Assume that the activation $\sigma\in C(\mathbb{R})$ satisfies the univariate universal approximation property of Definition 2.2. Then for every continuous functional $f\in C(V)$ and every $\varepsilon>0$ there exist integers $N,k\geq 1$, points $x_{1},\dots,x_{k}\in K$, and real parameters $c_{i},\theta_{i},\xi_{ij}$ ($i=1,\dots,N$, $j=1,\dots,k$) such that
@@
We view $C(Y)$ as a Banach space and hence as a Hausdorff locally convex space. Apply Theorem 2.2 with $X=C(Y)$, compact set $V\subset X$, and $m=1$. Then there exist $N\geq 1$, functionals $\ell_{1},\dots,\ell_{N}\in C(Y)^{*}$, and parameters $c_{i},\theta_{i}\in\mathbb{R}$ such that
@@
Theorem 2.2 shows that locally convex spaces (including Banach spaces and Fréchet spaces) form a broad and natural class of input domains for which universality of topological feedforward neural networks holds. This universality result is driven by the abstract $D$-property and relies on the separating power of the continuous dual, rather than on any finite-dimensional structure of the input space. The subsequent Theorem 2.3 of Chen and Chen may be viewed as a special case of this general principle applied to the Banach space $C(Y)$. In the next section, we turn to the substantially more delicate question of whether universality can be retained under uniform width constraints, that is, in the deep narrow framework.
@@ 3 Universality of deep narrow TFNNs
In this section we address the question announced in the previous remark: whether universal approximation on a general topological space $X$ can be achieved by deep networks of uniformly bounded width. As discussed earlier, this problem is more delicate than universality without width constraints. The $D$-property yields approximation by sums of many univariate compositions, but it does not by itself control how many neurons must appear in each hidden layer.
Our strategy is to impose an additional finite-dimensional composition assumption on the feature family $\mathcal{A}(X)$. Under this assumption, approximation on $X$ reduces to approximation on a compact subset of $\mathbb{R}^{n}$. We can then invoke the deep narrow universality theorem of Kidger and Lyons [18] for Euclidean inputs and vector-valued outputs.
@@ 3.1 Finite-dimensional composition via feature maps
3.2 Deep narrow TFNNs and a width constraint
We now introduce the width-constrained class that will be used in the sequel.
@@ 3.4 Deep narrow universality based on Ostrand inner functions
The present paper extends this viewpoint in two directions. First, we treat vector-valued outputs and formulate universality within a unified framework encompassing both shallow and deep architectures. Second, we address the width-constrained setting: deep narrow universality is established under a finite-dimensional composition property, which identifies when approximation on a general topological space can be reduced to approximation on a Euclidean space. In particular, for products of compact metric spaces we obtain explicit width bounds expressed in terms of topological dimension, thereby linking geometric and topological structure to quantitative architectural constraints.
@@
In contrast, in our framework the feature map $F$ is intrinsic to the neural network architecture on $X$ and is constructed from an admissible feature family $\mathcal{A}(X)\subset C(X)$. In the universality results without width constraints, the maps $F=(f_{1},\dots,f_{n})$ entering the first hidden layer of the network representation depend on the compact set under consideration, the target function, and the required accuracy, similarly to trainable parameters in classical networks. In the deep narrow setting, a finite-dimensional composition is imposed, and $F=(f_{1},\dots,f_{n})$ depends only on the compact set $K$ and the output dimension $m$. Thus, in both the unconstrained and deep narrow settings, the feature map $F$ is variable rather than a single fixed global injective map as in [20].
@@ 4 Discussion and Outlook
This work develops a general approximation-theoretic framework for neural networks whose inputs range over arbitrary topological spaces. By formulating neural architectures in terms of admissible feature families, we obtain a unified setting that encompasses classical feedforward networks on Euclidean spaces as a special case.
A central focus of the paper is the universal approximation property for both shallow and deep networks with vector-valued outputs. In the absence of width constraints, we establish general density results that extend classical universality theorems to non-Euclidean spaces.
The novelty also arises in the analysis of deep narrow networks, where the width of each hidden layer is uniformly bounded while depth is allowed to grow. We identify conditions under which such width-constrained networks retain universal approximation power, thereby extending deep narrow universality beyond the Euclidean setting.
An important outcome of this analysis is the explicit connection between universality of deep narrow networks and classical results from topology and dimension theory. In particular, Ostrand’s extension of the Kolmogorov superposition theorem yields concrete feature maps for products of compact metric spaces. As a consequence, we obtain explicit width bounds expressed in terms of topological dimension. This demonstrates how geometric and topological structure of the input space can be translated into approximation properties of the considered networks.
Several directions for further research remain open. One natural extension is the study of neural networks with outputs in infinite-dimensional spaces. Another promising direction is the investigation of quantitative approximation rates and complexity bounds within the present topological framework. Finally, it would be of interest to explore how the abstract feature-based approach developed here interfaces with practical applications.

### v2

@@ 2.4 Locally convex spaces
Let $X$ be a locally convex topological vector space and let $\sigma:\mathbb{R}\to\mathbb{R}$ satisfy the univariate universal approximation property of Definition 2.2. Then, for every integer $m\geq 1$, the TFNN class $\mathcal{N}_{1}^{(m)}(\sigma)$ associated with the feature family $\mathcal{A}(X)=X^{*}$ is dense in $C(X;\mathbb{R}^{m})$ with respect to the topology of uniform convergence on compact sets.
@@
Moreover, $\mathcal{E}$ separates points of $K$. Indeed, since $X$ is locally convex, the Hahn–Banach continuous extension theorem holds. It is a consequence of this theorem that for any $x,y\in K$ with $x\neq y$ there exists $f\in X^{*}$ such that $f(x)\neq f(y)$ (see, e.g., [31, Theorem 3.6]). Consequently,
@@
We now apply Theorem 2.1. Let $m\geq 1$, let $K\subset X$ be compact, and let $g\in C(X;\mathbb{R}^{m})$ be arbitrary. Since $\mathcal{A}(X)=X^{*}$ has the $D$-property and $\sigma$ satisfies the univariate universal approximation property, Theorem 2.1 implies that for every $\varepsilon>0$ there exists a TFNN
@@
The theorem is proved. ∎
@@
Let $Y$ be a compact metric space and let $V\subset C(Y)$ be compact. Assume that the activation $\sigma\in C(\mathbb{R})$ satisfies the univariate universal approximation property of Definition 2.2. Then for every continuous functional $f\in C(V)$ and every $\varepsilon>0$ there exist integers $N,k\geq 1$, points $x_{1},\dots,x_{k}\in Y$, and real parameters $c_{i},\theta_{i},\xi_{ij}$ ($i=1,\dots,N$, $j=1,\dots,k$) such that
@@
We view $C(Y)$ as a Banach space and hence as a Hausdorff locally convex space. Since $C(Y)$ is a metric space, it is normal, and the compact set $V\subset C(Y)$ is closed. Therefore, by the Tietze extension theorem, the function $f\in C(V)$ admits a continuous extension $F\in C(C(Y))$. Applying Theorem 2.2 to $F$ with $X=C(Y)$ and $m=1$, we obtain $N\geq 1$, functionals $\ell_{1},\dots,\ell_{N}\in C(Y)^{*}$, and parameters $c_{i},\theta_{i}\in\mathbb{R}$ such that
$\sup_{u\in V}\left|F(u)-\sum_{i=1}^{N}c_{i}\,\sigma(\ell_{i}(u)-\theta_{i})\right|<\frac{\varepsilon}{2}.$
Since $F=f$ on $V$, it follows that
@@
Theorem 2.2 shows that locally convex spaces (including Banach spaces and Fréchet spaces) form a broad and natural class of input domains for which universality of topological feedforward neural networks holds. This universality result is based on the abstract $D$-property and the separating power of the continuous dual. The subsequent Theorem 2.3 of Chen and Chen may be viewed as a special case of this general principle applied to the Banach space $C(Y)$. In the next section, we turn to the substantially more delicate question of whether universality can be retained under uniform width constraints, that is, in the deep narrow framework.
@@ 3 Universality of deep narrow TFNNs
In this section we study whether universal approximation on a general topological space $X$ can be achieved by deep neural networks with uniformly bounded width. To address this question, it is not sufficient to rely solely on the $D$-property. While this property ensures approximation by finite sums of univariate compositions, it does not provide control over the number of neurons required in each hidden layer.
To overcome this limitation, we impose a finite-dimensional composition assumption on the feature family $\mathcal{A}(X)$. Under this assumption, approximation on $X$ can be reduced to approximation on a compact subset of $\mathbb{R}^{n}$, where known deep narrow universality results, such as those of Gripenberg [8] and Kidger and Lyons [18], apply.
@@ 3.1 Finite-dimensional composition via feature maps
3.2 Deep narrow TFNNs
We now consider the class of bounded-width deep networks that will be used in the sequel.
@@ 3.4 Deep narrow universality based on Ostrand inner functions
The present paper extends this viewpoint in two directions. First, we treat vector-valued outputs and consider both shallow and deep architectures. Second, we address the bounded-width setting: deep narrow universality is established under a finite-dimensional composition property, which identifies when approximation on a general topological space can be reduced to approximation on a Euclidean space. In particular, for products of compact metric spaces we obtain explicit width bounds expressed in terms of topological dimension, thereby linking geometric and topological structure to quantitative architectural constraints.
@@
In contrast, in our framework the feature map $F$ is intrinsic to the neural network architecture on $X$ and is constructed from an admissible feature family $\mathcal{A}(X)\subset C(X)$. In the universality results without width constraints, the maps $F=(f_{1},\dots,f_{n})$ entering the first hidden layer of the network representation depend on the compact set under consideration, the target function, and the required accuracy, similarly to trainable parameters in classical networks. In the deep narrow setting, a finite-dimensional composition is imposed, and $F=(f_{1},\dots,f_{n})$ depends only on the compact set $K$ and the output dimension $m$. Thus, in both the arbitrary-width and bounded-width settings, the feature map $F$ is variable rather than a single fixed global injective map as in [20].
@@ 4 Discussion and Outlook
This paper develops a general approximation-theoretic framework for neural networks whose inputs range over arbitrary topological spaces. By formulating neural architectures in terms of admissible feature families, we obtain a unified setting that encompasses classical feedforward networks on Euclidean spaces as a special case.
We study the universal approximation property for topological neural networks with vector-valued outputs. In the arbitrary-width case, we establish general density results that extend classical universality theorems to non-Euclidean spaces.
We also analyze deep narrow networks, where the width of each hidden layer is uniformly bounded while depth is allowed to grow. We identify conditions under which such networks retain the universal approximation property, thereby extending deep narrow universality beyond the Euclidean setting.
A consequence of this analysis is the explicit connection between universality of deep narrow networks and classical results from topology and dimension theory. In particular, Ostrand’s extension of the Kolmogorov superposition theorem yields concrete feature maps for products of compact metric spaces. As a consequence, we obtain explicit width bounds expressed in terms of topological dimension. This demonstrates how geometric and topological structure of the input space can be translated into approximation properties of the considered networks.
Several directions for further research remain open. One natural extension is the study of neural networks with outputs in infinite-dimensional spaces. Another direction is the investigation of quantitative approximation rates and complexity bounds within the present topological framework. Finally, it would be of interest to explore how the abstract feature-based approach developed here interacts with practical applications.